Review (closed economy)

C=C0 +0*8 yd

I=150-10i

G=250

T=200

Ms=100

Md=40+0.ıy -10i

A) Write the equation for IS

B)Write the equation for LM

C) Find the equilibrium valuer for income, Y, and interest rate, İ, (IS/LM)

D)Suppose that, investment is completely interest inelastic , Write new IS-LM equations?

E) Find new equilibrium values for income and interest rates

F) Draw IS-LM diagrams


1
Expert's answer
2022-03-27T18:52:07-0400

a. The condition of t he equilibrium of the market of goods in the Keynesian model: the income of the economic system at the point of equilibrium is equal to expenses. We find the equation of the line IS, which determines all the points of equilibrium of the market of goods.

"Y=C+I+G"

"Y=60 + 0.8\times YD+150-10\times r+250"

"Y=60 + 0.8(Y-T)+150-10r+250"

"Y=60 + 0.8(Y-200)+150-10r+250"

"Y=300+ 0.8\times Y-10r"

"Y-0.8\times Y=300-10\times r"

"0.2\times Y=300-10\times r"

"Y=1500-50\times r" the IS equation

Ms=Md

"100=40 + 0.1\times Y - 10\times r"

"60= 0.1\times Y - 10\times r"

"600+100\times r=Y" the LM equation


b.IS=LM

"1500-50\times r=600+100\times r"

r=6%


"Y=1500-50\times r=1500-50\times6=1500-300=1200"